OpenStudy (anonymous):

4 − (x + 2) = 6 − 5(x + 4) Can someone please help me find the value of x?

OpenStudy (anonymous):

ok 4-(x+2)=6-5(x+4) first you need to simplyfy and combine like terms so -4x+8=6-5x+-20 combine like terms -9x+-12 (I think I did something wrong) but now d0 9/12 and that will solve x, but if its supposed to be a nondecimal then maybe simplyfy a different way

OpenStudy (jack1):

4 − (x + 2) = 6 − 5(x + 4) 4 - 2 - x = 6 - 5x - 20 2 - x = -14 - 5x -x + 5x = -14 - 2 4x = -16 so x =...@Jay5 ?

OpenStudy (jack1):

@BeckieBird ur only error wat u times the LHS by 4 when it was 4 - not 4 times, otherwise would have been perfect

OpenStudy (anonymous):

aw ok. thanks :D

OpenStudy (jack1):

np hay, we pals now ;D @Jay5 did u catch all or the above and understand?

OpenStudy (anonymous):

Yea. Thanks(: